Walt made an extra $9000 last year from a part-time job. He invested part of the money at a 9% annual rate of interest, and the rest at an 8% annual rate of interest. At the end of the year, he made a total of $770 in interest. How much was invested at 8%?

Answer by stanbon(75887)

Walt made an extra $9000 last year from a part-time job. He invested part of the money at a 9% annual rate of interest, and the rest at an 8% annual rate of interest. At the end of the year, he made a total of $770 in interest. How much was invested at 8%?
--------------
n + e = 9000
0.09n + 0.08e = 770
--------------------------
Multiply thru 1st eq. by 9
Multiply thru 2nd eq. by 100
------
9n + 9e = 9*9000
9n + 8e = 77000
----------------------
Subtract and solve for "e":
e = 4000 (amt invested at eight%)
======================================
Cheers,
stan H.
============